OpenStudy (anonymous):

Solve for (0, -2), (3,2) in Slope-Intercept Form!

OpenStudy (nubeer):

first u have to find slope m = (y2-y1)/(x2-x1)

OpenStudy (nubeer):

after u find slope pick any of the point given i will pick (0,-2) x1= 0, y1 = -2 now use this equation y-y1 = m(x-x1) plug in the values

OpenStudy (anonymous):

thanks